# Retrieval-Augmented Decision Transformer: External Memory for In-context RL

### v2

@@ 3.2.2 Searching for Similar Experiences
where $\operatorname{cossim}(\bm{q},\bm{k})=\frac{\mathbf{q}\cdot\mathbf{k}}{\|\mathbf{q}\|\|\mathbf{k}\|}$ is the cosine similarity. Consequently, $\mathcal{R}$ contains the set of retrieved sub-trajectories and their keys. Providing too similar experiences to the model may hinder learning [Yasunaga et al., 2023] and we apply retrieval regularization during training (see Appendix C.4).
@@ 3.2.3 Reweighting Retrieved Experiences
Following Park et al. [2023], we characterize the usefulness of retrieved sub-trajectories in $\mathcal{R}$ along two dimensions: relevance and utility. The relevance of a key $\bm{k}\in\mathcal{K}$ is defined by its cosine similarity to the query $\bm{q}$. While a retrieved experience may be relevant, it might not be important. Determining the utility of a sequence in general is hard. Thus, we experiment with two heuristics that follow different definitions of utility. The first assigns more utility to sub-trajectories with high return, and is utilized at inference only. The second assigns utility to sub-trajectories that originate from the same task as the query and is used at training only. Then, we reweight a retrieved experience according to:
@@
During inference, we evaluate the ICL capabilities of the agent. Starting from an empty vector index, we store experiences of the agent while it interacts with the environment (see Figure 1, right). Thus, during inference, the agent can only retrieve experiences from the same task. Therefore, we steer the agent to produce high reward behaviour on the new task by reweighting a retrieved sub-trajectory by the total return achieved over the episode it appears in, i.e., $s_{\text{u}}(\tau_{\text{ret}},\tau_{\text{in}})=\sum_{i=0}^{T}r_{i}$. We apply this reweighting to the retrieved experiences in $\mathcal{R}$ and select the top-$k$ elements by:
@@
0:  DT $\pi_{\theta}$, embed model $g$, episodes $N$, episode len $T$, context len $C$, retrieve, reweight.
1:  $\mathcal{I}\leftarrow\emptyset$triangleright Inititalize index
@@
5:        $\bm{q}=g(\tau_{t-C:t})$triangleright Construct query
6:        $\mathcal{R}\leftarrow\text{{retrieve}}(\bm{q},\mathcal{I})$triangleright Top-$l$ trjs, Eq. 3
7:        $\mathcal{S}\leftarrow\text{{reweight}}(\mathcal{R})$triangleright Top-$k$, Eq. 4, 5
8:        $a\sim\pi_{\theta}(a\mid\tau_{t-C:t},\{\tau_{\text{ret}}\in\mathcal{S}\})$triangleright Predict
@@
10:        $\tau\leftarrow\tau\cup(s,a,r)$triangleright Append transition to $\tau$
@@ 3.2.4 Incorporating Retrieved
(a) Dark-Room 10$\times$10
(b) Dark-Room 20$\times$20
(c) Dark-Room 40$\times$20
Figure 3: ICL performance on Dark-Room (a) 10$\times$10, (b) 20$\times$20, (c) 40$\times$20 at end of training (100K steps). We evaluate each agent for 40 episodes on each of the 20 evaluation tasks and report mean reward (+ 95% CI) over 3 seeds.
@@ 4 Experiments
We evaluate the ICL abilities of RA-DT on grid-world environments used in prior works, namely Dark-Room (see Section 4.1), Dark Key-Door (Section 4.2), and MazeRunner (Section 4.3) [Laskin et al., 2022; Lee et al., 2022; Grigsby et al., 2023], with increasingly larger grid-sizes, resulting in longer episodes. Moreover, we evaluate RA-DT on two robotic benchmarks (Meta-World and DMControl, Section 4.4) and procedurally-generated video games (Procgen, Section 4.5).
Across experiments, we report performances for two variants of RA-DT. The first variant leverages a domain-specific embedding model for retrieval, specifically a DT trained on the same domain. The second variant (RA-DT + Domain-agnostic) makes use of the FH mechanism in combination with BERT [Devlin et al., 2019] as the pre-trained LM. Consequently, this variant of RA-DT does not require any domain-specific pre-training of the embedding model. We compare RA-DT against the vanilla DT and two established in-context RL methods, namely Algorithm Distillation [Laskin et al., 2022, AD] and Decision Pre-trained Transformer [Lee et al., 2023, DPT]. Following, Agarwal et al. [2021] we report the mean across tasks and 95% confidence intervals over 3 seeds. We use a context length equivalent to two episodes (from 200 up to 2000 timesteps) for AD, DPT and DT. For RA-DT, we use a considerably shorter context length of 50 transitions, unless mentioned otherwise. On grid-worlds, we train all methods for 100K steps and evaluate after every 25K steps. Similarly, we train for 200K steps and evaluate after every 50K steps for Meta-World, DMControl and Procgen. All grid-worlds and Procgen exhibit discrete actions and consequently, we train all methods via the cross-entropy loss to predict the next actions. On Meta-World and DMControl, we train all method using the mean-squared error loss to predict continuous actions. Following Laskin et al. [2022] and Lee et al. [2023], our primary evaluation criterion is performance improvement during ICL trials. After training, the agent interacts with the environment for a fixed amount of episodes, each of which is considered a single trial. Upon completion of an ICL trial, the respective episode is stored in the vector index. We provide further training and implementation details in Appendix C.
@@ 4.1 Dark-Room
Experiment Setup. Dark-Room is commonly used in prior work on in-context RL [Laskin et al., 2022; Lee et al., 2023]. The agent is located in an empty room, observes only its x-y coordinates, and has to navigate to an invisible goal state ($|\mathcal{S}|=2$, $|\mathcal{A}|=5$, see Figure 9). A reward of +1 is obtained in every step the agent is located in the goal state. Because of partial observability, it must leverage memory of previous episodes to find the goal. We conduct experiments on three different grid sizes, namely 10$\times$10, 20$\times$20, and 40$\times$20, and corresponding episode lengths of 100, 200 and 800, respectively. We designate 80 and 20 randomly assigned goals as train and evaluation locations, respectively, as in Lee et al. [2023]. We use Proximal Policy Optimization (PPO) [Schulman et al., 2017] to generate 100K transitions per goal for 10$\times$10 and 20$\times$20 grids and 200K for 40$\times$20 (see Figure 7 for single task expert scores). During evaluation, the agent interacts with the environment for 40 ICL trials, and we report the scores at the last evaluation step (100K). We provide additional details on the environment, the generated data, and the training procedure in Appendix B.1 and C.
Results. In Figure 3, we show the ICL performances on the 20 hold-out tasks for all considered methods on Dark-Room (a)10$\times$10, (b) 20$\times$20, and (c) 40$\times$20. In addition, we present the ICL curves on the training tasks and the learning curves across the entire training period in Figures 14 and 15 in Appendix D.1. Overall, we observe that RA-DT attains the highest average rewards on all 3 grid-sizes at the end of the 40 ICL-trials. On 10$\times$10, RA-DT obtains near-optimal performance scores both with the domain-specific and domain-agnostic embedding model. The vanilla DT does not exhibit any performance improvement across trials. This indicates the improvement in performance for RA-DT can be attributed to the retrieval component. Furthermore, RA-DT outperforms AD and DPT without keeping entire episodes in its context window. Similarly, RA-DT outperforms all baselines on the 20$\times$20 and 40$\times$20 grids. While RA-DT successfully improves in-context, the baselines exhibit only little learning progress over the ICL trials, especially for larger grid sizes. However, the final performance scores for 20$\times$20 and 40$\times$20 are not optimal. With increasing grid size, discovering the goal requires systematic exploration in combination with targeted exploitation. Therefore, we conduct a qualitative analysis on the exploration behaviour of RA-DT. We find that RA-DT develops strategies to imitate a given successful context (see Figure 16), and avoids low-reward routes given an unsuccessful context (see Figure 17).
(a) Dark Key-Door 10$\times$10
(b) Dark Key-Door 20$\times$20
(c) Dark Key-Door 40$\times$20
Figure 4: ICL performance on Dark Key-Door (a) 10$\times$10, (b) 20$\times$20, (c) 40$\times$20 at end of training (100K steps). We evaluate each agent for 40 episodes on each of the 20 evaluation tasks and report mean reward (+ 95% CI) over 3 seeds.
@@ 4.2 Dark Key-Door
Experiment Setup. In Dark Key-Door, the agent is located in a room with two invisible objects: a key and a door. The agent has to pick up the invisible key, then navigate to the door. Because of the presence of two key events, the task-space is combinatorial in the number of grid-cells ($10000100^{2}=$ 10000 possible tasks for $10\times 10$) and is therefore considered more difficult. A reward of +1 is obtained once for picking up the key and for every step the agent stands on the door grid-cell after it collected the key. We retain the same experiment setup as in Section 4.1 and provide further details in Appendix B.1 (also see Figure 8 for single-task expert scores).
@@ 4.3 Maze-Runner
Experiment Setup. Maze-Runner was introduced by Grigsby et al. [2023] and inspired by Pasukonis et al. [2022]. The agent is located in a procedurally-generated $15\times 15$ maze (see Figure 10), observes continuous Lidar-like depth representations of states, and has to navigate to one, two, or three goal locations in the correct order ($|\mathcal{S}|=6$,$|\mathcal{A}|=4$). A reward of +1 is obtained when reaching a goal location. Episodes last for a maximum of 400 steps, or terminate early if all goal locations have been visited. Similar to Dark-Room, we use PPO to generate 100K environment interactions for 100 procedurally-generated mazes. We train all methods on a multi-task dataset that comprises trajectories from 100 mazes, evaluate on 20 unseen mazes, and report performance over 30 ICL trials. We give further details on the environment, the dataset, and the experiment setup in Appendix B.2 and D.2.
@@ 4.4 Meta-World & DMControl
Experiment Setup. Next, we evaluate RA-DT on two multi-task robotics benchmarks, Meta-World [Yu et al., 2020b] and DMControl [Tassa et al., 2018]. States and actions in both benchmarks are multidimensional continuous vectors. While the state and action space in Meta-World remain constant across all tasks ($|\mathcal{S}|=39$, $|\mathcal{A}|=6$), they vary considerably in DMControl ($3\leq|\mathcal{S}|\leq 24$, $1\leq|\mathcal{A}|\leq 6$). Episodes last for 200 and 1000 steps in Meta-World and DMControl, respectively.
We leverage the datasets released by Schmied et al. [2024]. For Meta-World, we pre-train a multi-task policy on 45 of the 50 tasks (ML45, 90M transitions in total) and evaluate on the 5 remaining tasks (ML5). Similarly, on DMControl, we pre-train on 11 tasks (DMC11, 11M transitions in total) and evaluate on 5 unseen tasks (DMC5). We provide further details on the environments, datasets, and experiment setup in Appendices B.3 and D.3, and B.4 and D.4 for Meta-World and DMControl, respectively.
@@ 4.5 Procgen
Experiment Setup. Finally, we conduct experiments on Procgen [Cobbe et al., 2020], a benchmark consisting of 16 procedurally-generated video games, designed to test the generalization abilities of RL agents. The procedural generation in Procgen is controlled by setting an environment seed, which results in visually diverse observations for the same underlying task (see starpilot-example in Figure 12). In Procgen, the agent receives image-based inputs ($|\mathcal{S}|=$3$\times$64$\times$64). All 16 tasks share a discrete action space ($|\mathcal{A}|=15$). Rewards are either dense or sparse depending on the environment.
We follow Raparthy et al. [2023] and use 12 tasks for training (PG12) and 4 tasks for evaluation (PG4). First, we generate datasets by training task-specific PPO agents for 25M timesteps on 200 environment seeds per task in easy difficulty. Then, we pre-train a multi-task policy on the PG12 datasets (24M transitions in total, 2M per task). We leverage the procedural generation of Procgen and evaluate all models in three settings: training tasks - seen (PG12-Seen), training tasks - unseen (PG12-Unseen), and evaluation tasks - unseen (PG4). Additional details on the generated datasets and our environment setup are available in Appendices B.5 and D.5.
Results. Similar to our results on Meta-World and DMControl, we find that RA-DT improves average performance scores across all three settings compared to the baselines (see Figure 26 and Tables 5, 6, 7 in Appendix D.5), but no method exhibits in-context improvement during evaluation (Figure 27). We further discuss our negative results on Procgen, Meta-World, and DMControl in Section 5.
@@ 4.6 Ablations
To better understand the effect of learning with retrieval, we present a number of ablation studies on essential components in RA-DT conducted on Dark-Room $10\times 10$ (more details in Appendix E).
Retrieval outperforms sampling of experiences. To investigate the effect of learning with retrieved context, we substitute retrieval with random sampling, either over all tasks, or from the same task (see Figure 6a). We find that training with retrieval outperforms both sampling variants, highlighting the benefit of training with retrieval to improve ICL abilities. We hypothesise this is because retrieval constructs bursty sequences, which was found to be important for ICL [Chan et al., 2022].
@@
Figure 6: Ablations on important components in RA-DT conducted on Dark-Room 10$\times$10. We show (a) the effect of training with retrieval vs. sampling, (b) a sensitivity analysis on $\alpha$ as used in the re-weighting mechanism during training, and (c) the effect of leveraging different LMs as pre-trained embedding models for domain-agnostic retrieval.
@@
Effect of Retrieval Regularization. We evaluate with three retrieval regularization strategies to mitigate the effect of copying the context: deduplication, similarity cut-off, and query dropout. To evaluate their impact on ICL performance, we systematically removed each one from RA-DT (see Figure 30). We found the combination of all three to be effective and add them to our pipeline.
Different LMs for domain-agnostic RA-DT. Finally, we investigate how strongly domain-agnostic RA-DT is influenced by the choice of pre-trained LM for the embedding model. We compare our default choice BERT against other smaller/larger LMs (see Figure 36). We found that BERT performs best and performance decreases with smaller models.
Effect of Retrieval on Training/Inference Efficiency. Retrieval-augmentation adds computational overhead to the training/inference pipeline due to the cost of embedding the query and searching for similar experiences. However, we find that RA-DT results in significantly faster training times because of shorter context length (up to 7×7$\times$7 × see Appendix E.7). At inference-time, RA-DT is slightly slower compared to baselines when retrieving at every step, but exhibits similar inference speeds when retrieving less frequently (see Appendix E.8). Importantly, the retrieval mechanism in RA-DT enables access to the entirety of the experiences collected across all ICL trials with small additional cost. The ability to access a broader range of experiences may be a reason for its enhanced performance.
For additional ablations on RA-DT and on our baselines, we refer to Appendix E.
@@ 5 Discussion
Memory-Exploitation vs. Meta-learning Abilities. Current offline in-context RL methods are predominantly evaluated on contextual bandits or grid-worlds, such as Dark-Room [Laskin et al., 2022; Lee et al., 2023; Lin et al., 2023; Zisman et al., 2023; Sinii et al., 2023; Huang et al., 2024], which can only be solved by leveraging the context. However, it remains unclear to what extent the agent learns to learn in-context or simply copies from its context. Further, in our experiments on fully-observable environments (MetaWorld, DMControl, and Procgen), we did not observe ICL behaviour (see Appendices D.3, D.4, D.5). Therefore, it is necessary that future research on in-context RL disentangles the effects of memory and meta-learning abilities, similar to memory and credit-assignment [Ni et al., 2024], potentially on recent benchmarks [Nikulin et al., 2023; 2024].
@@
Challenges of Next-Action Prediction. Most in-context RL methods learn from offline datasets via next-action prediction and causal sequence modelling objectives. As such, they cannot learn to infer the utility of an action, and thus, distinguish between positive and negative examples. This can induce delusions, which lead to repetitions of suboptimal actions and copying behaviour [Ortega et al., 2021] (see Figure 19 for examples on Dark-Room). In contrast, online in-context RL methods have shown promising adaptation abilities [Team et al., 2023; Grigsby et al., 2023; Lu et al., 2024]. A similar trend has been observed in online meta-RL methods [Melo, 2022; Shala et al., 2024]. Consequently, a potential remedy to this problem is to train a value function to learn the utility of an action. However, this is usually not straightforward and requires constrained optimization objectives [Zanette et al., 2021; Kumar et al., 2020]. Therefore we leave this approach to future work.
@@
Conditioning Strategies in RL. In LLMs, applying sophisticated conditioning strategies is important to improve ICL abilities [Wei et al., 2022; Yao et al., 2024; Agarwal et al., 2024]. Even though RTG-conditioning [Chen et al., 2021], and chain-of-hindsight [Liu & Abbeel, 2023] have shown promise for generating high reward behaviour in DTs, the broader landscape for conditioning strategies for in-context RL remains under-explored. Therefore, we believe that systematically investigating conditioning methods for in-context RL is a fruitful direction for future research.
@@
Diversity of the Pre-training Distribution. The diversity and scale of the pre-training dataset may significantly affect the emergence of ICL. In our experiments, we pre-train on a relatively small set of tasks. Our results on gridworlds suggest that this is sufficient for ICL to emerge on simple environments. However, on more complex environments, the unseen tasks can be considered out-of-distribution and higher pre-training diversity may be necessary for ICL to emerge. It remains unclear how much diversity is required to elicit in-context RL, and if existing large-scale agents exhibit ICL [Reed et al., 2022; Raad et al., 2024]. One promising approach is to expand the pre-training diversity through learned interactive simulations [Yang et al., 2023a; Bruce et al., 2024].

### v3

@@ 3.2.2 Searching for Similar Experiences
where $\operatorname{cossim}(\bm{q},\bm{k})=\frac{\mathbf{q}\cdot\mathbf{k}}{\|\mathbf{q}\|\|\mathbf{k}\|}$ is the cosine similarity. Consequently, $\mathcal{R}$ contains the set of retrieved sub-trajectories and their keys. Providing too similar experiences to the model may hinder learning (Yasunaga et al., 2023), and we apply retrieval regularization during training (see Appendix C.4).
@@ 3.2.3 Reweighting Retrieved Experiences
Following Park et al. (2023), we characterize the usefulness of retrieved sub-trajectories in $\mathcal{R}$ along two dimensions: relevance and utility. The relevance of a key $\bm{k}\in\mathcal{K}$ is defined by its cosine similarity to the query $\bm{q}$. While a retrieved experience may be relevant, it might not be important. Determining the utility of a sequence in general is difficult. Therefore, we introduce two heuristics that follow different definitions of utility. The first assigns more utility to sub-trajectories with high return, and is utilized at inference only. The second assigns utility to sub-trajectories that originate from the same task as the query and is used at training only. Then, we reweight a retrieved experience according to:
@@
During inference, we evaluate the ICL capabilities of the agent. Starting from an empty vector index, we store experiences of the agent while it interacts with the environment (see Figure 1, right). During inference, the agent can only retrieve experiences from the same task. Consequently, we steer the agent to produce high reward behaviour on the new task by reweighting a retrieved sub-trajectory by the total return achieved over the episode it appears in, i.e., $s_{\text{u}}(\tau_{\text{ret}},\tau_{\text{in}})=\sum_{i=0}^{T}r_{i}$. We apply this reweighting to the retrieved experiences in $\mathcal{R}$ and select the top-$k$ elements by:
@@
0: DT $\pi_{\theta}$, embed model $g$, episodes NN, episode len $T$, context len $C$, retrieve, reweight.
[…]
@@
5:   $\bm{q}=g(\tau_{t-C:t})$ $\triangleright$ Construct query
6:   $\mathcal{R}\leftarrow\text{{retrieve}}(\bm{q},\mathcal{I})$ $\triangleright$ Top-$l$ trjs, Eq. 3
7:   $\mathcal{S}\leftarrow\text{{reweight}}(\mathcal{R})$ $\triangleright$ Top-$k$, Eq. 4, 5
8:   $a\sim\pi_{\theta}(a\mid\tau_{t-C:t},\{\tau_{\text{ret}}\in\mathcal{S}\})$ $\triangleright$ Predict
@@
10:   $\tau\leftarrow\tau\cup(s,a,r)$ $\triangleright$ Append transition to $\tau$
@@ 4 Experiments
We evaluate the ICL abilities of RA-DT on grid-world environments used in prior works, namely Dark-Room (see Section 4.1), Dark Key-Door (Section 4.2), and MazeRunner (Section 4.3) (Laskin et al., 2022; Lee et al., 2022; Grigsby et al., 2023), with increasingly larger grid-sizes, resulting in longer episodes. Moreover, we evaluate RA-DT on two robotic benchmarks (Meta-World and DMControl, Section 4.4) and procedurally-generated video games (Procgen, Section 4.5).
Across experiments, we report performances for two variants of RA-DT. The first variant leverages a domain-specific embedding model for retrieval, specifically a DT trained on the same domain. The second variant (RA-DT + Domain-agnostic) makes use of the FH mechanism in combination with BERT (Devlin et al., 2019) as the pre-trained LM. Consequently, this variant of RA-DT does not require any domain-specific pre-training of the embedding model. We compare RA-DT against the vanilla DT and two established in-context RL methods, namely Algorithm Distillation (Laskin et al., 2022, AD) and Decision Pre-trained Transformer (Lee et al., 2023, DPT). Following Agarwal et al. (2021), we report the mean across tasks and 95% confidence intervals over 3 seeds. We use a context length equivalent to two episodes (from 200 up to 2000 timesteps) for AD, DPT, and DT. For RA-DT, we use a considerably shorter context length of 50 transitions, unless mentioned otherwise. On grid-worlds, we train all methods for 100K steps and evaluate after every 25K steps. Similarly, we train for 200K steps and evaluate after every 50K steps for Meta-World, DMControl, and Procgen. All grid-worlds and Procgen exhibit discrete actions, and consequently, we train all methods via the cross-entropy loss to predict the next actions. On Meta-World and DMControl, we train all methods using the mean-squared error loss to predict continuous actions. Following Laskin et al. (2022) and Lee et al. (2023), our primary evaluation criterion is performance improvement during ICL trials. After training, the agent interacts with the environment for a fixed number of episodes, each of which is considered a single trial. Upon completion of an ICL trial, the respective episode is stored in the vector index. We provide further training/implementation details in Appendix C.
(a) Dark-Room $\times10$
(b) Dark-Room $\times20$
[…]
@@ 4.1 Dark-Room
[…]
(a) Dark Key-Door $\times10$
(b) Dark Key-Door $\times20$
[…]
@@ 4.2 Dark Key-Door
Experiment Setup. In Dark Key-Door, the agent is located in a room with two invisible objects: a key and a door. The agent has to pick up the invisible key, then navigate to the door. Because of the presence of two key events, the task-space is combinatorial in the number of grid-cells ($100^{2}=10000$ possible tasks for $10\times 10$) and is therefore considered more difficult. A reward of +1 is obtained once for picking up the key and for every step the agent stands on the door grid-cell after it has collected the key. We retain the same experiment setup as in Section 4.1 and provide further details in Appendix B.1 (also see Figure 8 for single-task expert scores).
@@ 4.3 Maze-Runner
Experiment Setup. Maze-Runner was introduced by Grigsby et al. (2023) and inspired by Pasukonis et al. (2022). The agent is located in a procedurally-generated $15\times 15$ maze (see Figure 10), observes continuous Lidar-like depth representations of states, and has to navigate to one, two, or three goal locations in the correct order ($|\mathcal{S}|=6$$|\mathcal{A}|=4$). A reward of +1 is obtained when reaching a goal location. Episodes last for a maximum of 400 steps, or terminate early if all goal locations have been visited. Similar to Dark-Room, we use PPO to generate 100K environment interactions for 100 procedurally-generated mazes. We train all methods on a multi-task dataset that comprises trajectories from 100 mazes, evaluate on 20 unseen mazes, and report performance over 30 ICL trials. We give further details on the environment/dataset/experiment setup in Appendix B.2 and D.2.
@@ 4.4 Meta-World & DMControl
Experiment Setup. Next, we evaluate RA-DT on two multi-task robotics benchmarks, Meta-World (Yu et al., 2020b) and DMControl (Tassa et al., 2018). States and actions in both benchmarks are multidimensional continuous vectors. While the state and action space in Meta-World remain constant across all tasks ($|\mathcal{S}|=39$, $|\mathcal{A}|=6$), they vary considerably in DMControl ($3\leq|\mathcal{S}|\leq 24$, $1\leq|\mathcal{A}|\leq 6$). Episodes last for 200 and 1000 steps in Meta-World and DMControl, respectively. We leverage the datasets released by Schmied et al. (2024). For Meta-World, we pre-train a multi-task policy on 45 of the 50 tasks (ML45, 90M transitions in total) and evaluate on the 5 remaining tasks (ML5). Similarly, on DMControl, we pre-train on 11 tasks (DMC11, 11M transitions in total) and evaluate on 5 unseen tasks (DMC5). We provide additional details on the environments/datasets/setup in Appendices B.3, D.3, B.4, and D.4.
@@ 4.5 Procgen
[…]
Results. Similar to Meta-World and DMControl, we find that RA-DT improves average performance scores across all three settings compared to the baselines (see Figure 26 and Tables 5, 6, 7 in Appendix D.5), but no method exhibits ICL during evaluation (Figure 27). We discuss our negative results on Procgen/Meta-World/DMControl in Section 5.
@@ 4.6 Ablations
To better understand the effect of learning with retrieval, we present a number of ablation studies on essential components in RA-DT conducted on Dark-Room $10\times 10$. We provide additional details on our ablations in Appendix E.
@@
(c) Training Efficiency, $40\times 20$
Figure 6: Ablations on important components in RA-DT. We show (a) the effect of training with retrieval vs. sampling, (b) a sensitivity analysis on $\alpha$ as used in the re-weighting mechanism during training on Dark-Room $\times10$. In (c), we show the training efficiency in terms of samples per second across methods on Dark-Room $40\times 20$
Retrieval outperforms sampling of experiences. To investigate the effect of learning with retrieved context, we substitute retrieval with random sampling, either over all tasks or from the same task (see Figure 6a). We find that training with retrieval outperforms both sampling variants, highlighting the benefit of training with retrieval to improve ICL abilities. This is because retrieval constructs bursty sequences, which is important for ICL (Chan et al., 2022).
@@
Different LMs for domain-agnostic RA-DT. We investigate how strongly domain-agnostic RA-DT is influenced by the choice of pre-trained LM for the embedding model. We compare our default choice, BERT, against smaller/larger LMs (see Figure 35). BERT performs best, and performance decreases with smaller models.
Effect of Retrieval on Training/Inference Efficiency. Retrieval-augmentation adds computational overhead to the training/inference pipeline due to embedding the query and searching for similar experiences. However, we find that RA-DT results in significantly faster training times because of shorter context length (see Appendix E.7). For the largest grid ($40\times 20$) and consequently episode length, we find that RA-DT is almost $7\times$ faster at training time (see Figure 6c). At inference time, RA-DT is slightly slower compared to baselines when retrieving at every step, but exhibits similar inference speeds when retrieving less frequently (see Appendix E.8). Furthermore, in Figure 34 we highlight the importance of FlashAttention (Dao, 2023) for handling long context. Importantly, the retrieval mechanism in RA-DT enables access to all collected experience with small additional cost, which is the reason for enhanced performance.
@@ 5 Discussion
Memory-Exploitation vs. Meta-learning Abilities. Current offline in-context RL methods are predominantly evaluated on contextual bandits or grid-worlds, such as Dark-Room (Laskin et al., 2022; Lee et al., 2023; Lin et al., 2023; Zisman et al., 2023; Sinii et al., 2023; Huang et al., 2024), which can only be solved by leveraging the context. It remains unclear to what extent the agent learns to learn in context or copies from its context. Further, in our experiments on fully-observable environments, we did not observe ICL behaviour (see Appendices D.3, D.4, D.5). Therefore, it is necessary that future research on in-context RL disentangles the effects of memory and meta-learning abilities, similar to memory and credit-assignment (Ni et al., 2024), potentially on recent benchmarks (Nikulin et al., 2023; 2024). We believe our datasets facilitate future work in this direction.
@@
Challenges of Next-Action Prediction. Most in-context RL methods learn from offline datasets via next-action prediction and causal sequence modelling objectives. As such, they cannot learn to infer the utility of an action, and thus, distinguish between positive/negative examples. This can induce delusions, which lead to repetitions of suboptimal actions and copying behaviour (Ortega et al., 2021) (see Figure 19 for examples on Dark-Room). In contrast, online in-context RL (Team et al., 2023; Grigsby et al., 2023; Lu et al., 2024) and online meta-RL methods (Melo, 2022; Shala et al., 2024) have shown promising adaptation abilities. A potential remedy to this problem is to train a value function to learn the utility of an action (Zanette et al., 2021; Kumar et al., 2020).
@@
Conditioning Strategies in RL. In LLMs, applying sophisticated conditioning strategies is important to improve ICL abilities (Wei et al., 2022; Yao et al., 2024; Agarwal et al., 2024). Finding appropriate conditioning strategies is also crucial for effective in-context RL. Even though RTG-conditioning (Chen et al., 2021) and chain-of-hindsight (Liu & Abbeel, 2023) have shown promise for generating high reward behaviour in DTs, the broader landscape for conditioning strategies for in-context RL remains under-explored. Therefore, we believe that systematically investigating conditioning methods for in-context RL is a fruitful direction for future research.
@@
Diversity of the Pre-training Distribution. The diversity and scale of the pre-training dataset significantly affect the emergence of ICL. In our experiments, we pre-train on a relatively small set of tasks. Our results on gridworlds suggest that this is sufficient for ICL to emerge in simple environments. In complex environments, unseen tasks can be considered out-of-distribution, and higher pre-training diversity may be necessary. It remains unclear how much diversity is required to elicit in-context RL, and if existing large-scale agents exhibit ICL (Reed et al., 2022; Raad et al., 2024). A promising approach is to expand diversity through learned interactive simulations (Bruce et al., 2024).